Hi,
Solve the sysytem of equations.
7X-2Y=3
3X+5Y=13
thank you!

Answer by Alan3354(21608)

7X-2Y=3
3X+5Y=13
-----------
7X-2Y-3=0
3X+5Y-13=0
Using determinants:
7 -2 -3
3 5 -13
x = 1
y = 2
It checks.
There are other ways to do it, but you can use Excel to do this with determinants.